An environmental economic consulting firm is hired to measure the negative externalities associated with the pollution from an industry. The consultants calculate the marginal social cost of production to be MSC 20+30 and the marginal private cost of production to be MPC-Q+30. The market demand curve can be expressed as P 60-Q. The sum of the consumer surplus plus producer surplus at the market equilibrium (i.e. the equilibrium where production decisions do not take into account the externality) is: A. $0. B. $225 O C. $112.5 D. $150

Steo by step explanations are appreciated Thank you
Answer #1

When externality doesn't take into account MC=MPC i.e Q+30

This represents the supply curve equation . When Q=0 , MPC=30.

Market demand , P=60-Q

When Q=0, P=60

At equilibrium P=MPC

60-Q= Q+30

2Q = 30

Q =15 units (equilibrium quantity)

P= 60-Q

P= 60-15 = $ 45 (Equilibrium price)

Consumer surplus is the triangle area above $45 and below the demand curve.

CS= (0.5)(60-45)(15)= $ 112.5

Producer surplus is the triangle area above the supply curve and below $45.

PS = (0.5)(45-30)(15) = $ 112.5

Total surplus = CS + PS = $ (112.5 +112.5)= $ 225

Hence, option(B) is correct.